OpenStudy (rva.raghav):

Disssolving 120g of urea(mol wt 60)in 1000g of water gave a sol of density 1.15g/ml.the molarity of the solution is ....? please tell me detailed steps.

OpenStudy (rva.raghav):

Please tell me soon

OpenStudy (anonymous):

Do u knw wat is Molarity..?

OpenStudy (anonymous):

\[D=\frac{ M }{ V }\] \[V=\frac{ M }{ D }\] M = 1000 + 120 = 1120gm

OpenStudy (rva.raghav):

but the answer is 2.05M

OpenStudy (anonymous):

Lol....First Find V then use Molarity = no of moles of solute / Volume (L)

OpenStudy (rva.raghav):

how to find V

OpenStudy (anonymous):

hm..thats wat i have Given u.....)

OpenStudy (anonymous):

1.15 g/ml = 1150 g/L

OpenStudy (anonymous):

\[V = 1120/1150 = 0.973L\] no of moles = 120/60 = 2

OpenStudy (anonymous):

\[M = 2/0.973 = 2.05M\]

OpenStudy (anonymous):

hope u got it...)

OpenStudy (rva.raghav):

tks

OpenStudy (anonymous):

Welcome..)

OpenStudy (rva.raghav):

yahoo...yahooo......oooooooooooooooo..............

OpenStudy (rva.raghav):

please wait i have to ask u another question

OpenStudy (rva.raghav):

Why NPN transistors a more preferred than PNP?